OpenStudy (anonymous):

Two similar parallelograms have areas of 18 in^2 and 32 in^2. Find the ratio of their two perimeters.

ganeshie8 (ganeshie8):

take the ratio of their areas, and take the square root

ganeshie8 (ganeshie8):

ratio of areas = \(\large \frac{18}{32} = \frac{9}{16}\)

ganeshie8 (ganeshie8):

now simply take the square root of above ratio

OpenStudy (anonymous):



ganeshie8 (ganeshie8):

noo its just 9/16

ganeshie8 (ganeshie8):

when u take square root, you would get 3/4 so ratio of perimeters = 3/4

ganeshie8 (ganeshie8):

getting ?

OpenStudy (anonymous):

yep

ganeshie8 (ganeshie8):

cool :)